Physics Drill: Ohm’s Law. Solve the problem using the principles of Ohm’s


Law below. Show your complete solution.


1. Find the current through a resistor of resistance 𝑅 = 2Ω if the voltage across


the resistor is 6V.


2. While repairing an electric bulb socket, an electrician gets a mild shock when a


current of 0.0050 A passes through him. The same electrician was killed by a


current of 0.15 A when he turned on an electric bulb while taking a bath. The


voltage in each situation is equal to 120 V. Find the resistance of the electrician


in each situation.


3. A typical value of the resistance of a person from hand to hand through the


body is 2,000 Ω. If a person with this resistance accidentally touches a 220 V


live wire, what current passes through the person?

1
Expert's answer
2022-03-27T18:21:35-0400

Answers


1.

"\qquad\qquad
\begin{aligned}
\small i&=\small \frac{V}{R}=\frac{6V}{2\Omega}
\end{aligned}"

2.

  • Apply the same formula used in the previous step. "\small R =\frac{V}{i}"


3.

  • Just apply the "\small V =iR" and get the current found.